# Parity violation as enforced symmetry breaking in 3D fermionic topological order

### v1

@@ Abstract
In the previous work Ref.[13], we systematically studied the ESB for 2D invertible fTO. Herein, we venture into the ESB phenomenon of 3D fTO. As there is no invertible fTO in 3D, we have to consider intrinsic fTO. In this work, we will focus on 3D ESB of TRS. We show that TRS is intrinsically broken by the presence of certain 3D fermionic TO (fTO) characterized by exotic non-Abelian three-loop braiding statistics. Such an fTO is closely related to an intrinsic interacting fermionic SPT (fSPT) order, which is specifically protected by symmetry $K_{f}=\mathbb{Z}_{2}^{f}\times\mathbb{Z}_{2}\times\mathbb{Z}_{8}$, characterized by non-Abelian three-loop braiding statistics [14]. This fTO is obtained by gauging $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ in the fSPT state, and will be referred to as Ising fTO. As we shall show, this $K_{f}$ fSPT is incompatible with $\mathcal{T}^{2}=1$ while compatible with $\mathcal{T}^{2}=\mathcal{P}_{f}$. Its incompatibility with the $\mathcal{T}^{2}=1$ TRS becomes exactly the ESB of TRS by this fTO. To our knowledge, this is the first and simplest example of TRS incompatibility of 3D fTO. This example readily extends to the realm of parity symmetry through the crystalline equivalence principle, while the concept of ESB unveils a novel pathway for comprehending parity violation within the Standard Model (SM), under the assumption that its vacuum acquires emergent generalized symmetry[15, 16, 17] characterized as certain 3D fTO. We will also discuss the general framework for 3D ESB phenomena.
@@
Upon gauging $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ of this fSPT (promoting the global $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ symmetry to be a local gauge symmetry), we obtain a 3D fTO, dubbed Ising fTO, with two elementary particle-like topological excitations: gauge charges $\mathsf{e}_{1,2}$, together with two elementary loop excitations $\Omega_{1,2}$, containing unit flux of $\mathbb{Z}_{2}$ and $\mathbb{Z}_{8}$ gauge group, respectively. Besides, there is also fermionic excitation that can be created by local operators and is usually called the transparent fermion $f$. General loop excitations contain the gauge flux and gauge charge. The flux of loop excitations forms the fusion group $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$. This fTO has many exotic properties. First, when $\Omega_{1}$ is linked to the base loop $\Omega_{2}$, its topological spin $\theta_{\Omega_{1};\Omega_{2}}$ satisfies $\Theta_{\Omega_{1};\leavevmode\nobreak\ \Omega_{2}}=2\theta_{\Omega_{1};\leavevmode\nobreak\ \Omega_{2}}=\frac{\pi}{4}$ mod $\frac{\pi}{2}$. Secondly, for two linked $\Omega_{1}$ and $\Omega_{2}$, each carries a Majorana zero mode (MZM), indicating the non-Abelian nature (see Fig.1c) [19, 14]. Thirdly, when further gauging the fermion parity, this resulting TO has fascinating statistics among loops, one of which is called the non-Abelian Ising three-loop braiding statistics as seen in Fig.1b . We note that in general the topological invariant $\Theta_{\Omega_{1};\leavevmode\nobreak\ \Omega_{2}}=\frac{x\pi}{4}$ mod $2\pi$, and it is non-Abelian when $x$ takes odd integer. In addition, there is another topological invariant $\Theta_{\Omega_{2};\leavevmode\nobreak\ \Omega_{1}}$ which can take 00 or $\pi$ characterizing an Abelian bosonic TO of $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ Dijkgraaf-Witten models. (For completeness, a concise overview of fSPT and associated three-loop braiding is provided in the supplementary materials). We also note that the ESB by TRS or parity symmetry we discuss below applies to these fTOs for all odd $x$.
@@
For three-loop statistics, this requirement leads to the conditions $\Theta_{\Omega_{i};\leavevmode\nobreak\ \Omega_{j}}=-\Theta_{\rho(\Omega_{i});\leavevmode\nobreak\ \rho(\Omega_{j})}$ with $i=1,j=2$ or $i=2,j=1$ where
@@
(see appendix for derivations). Solving the equations, we find (i) when $t=0$, $p=3$; (ii) when $t=1$, $p=3$ if $\Theta_{\Omega_{1},\Omega_{2};\leavevmode\nobreak\ \Omega_{2}}=0$ and $p=1$ if $\Theta_{\Omega_{1},\Omega_{2};\leavevmode\nobreak\ \Omega_{2}}=\pi$; (iii) $q=0$. Therefore, we are left with only three possibly consistent $\rho$, labeled by $\mathbf{r}=(3,0,0)$, $(3,0,1)$, and $(1,0,1)$, respectively.
@@
The structure of $G_{f}$ is described by: (1) how TRS acts on the subgroup $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$: $\rho_{\mathcal{T}}(\mathbf{g})=\mathbf{g}^{1+2p}\mathbf{h}^{q}$ and $\rho_{\mathcal{T}}(\mathbf{h})=\mathbf{g}^{4t}\mathbf{h}$, where $\mathbf{r}=(p,q,t)$ is the same as above [31]; and (2) what $\mathcal{T}^{2}$ is: $\mathcal{T}^{2}=\nu_{2}^{a}\nu_{2}^{b}\in K_{f}$, where $\nu_{2}^{a}=1$ or $\mathcal{P}_{f}$, and $\nu_{2}^{b}=1,\mathbf{g}^{2},\mathbf{g}^{4}$ for $\mathbf{r}=(3,0,0)$ or $(p,1,0)$ and $\nu_{2}^{b}=1$ for other choices of $\mathbf{r}$ (after considering some physical constraints for simplification  [^2] There are in total 28 different $G_{b}$, which will simplify to 17 cases after these physical constraints. However, we calculate all the 28 cases and found no case can reduce to the $K_{f}$ fSPT. ; see appendix). All 12 choices of $\mathbf{r}$ are considered, not limited to the three choices compatible with three-loop statistics. The Kramer-singlet and -doublet fermion systems correspond to $\nu_{2}^{a}=1$ and $P_{f}$, respectively. Below we demonstrate that ESB occurs for Kramer-singlet fermions. In this case, $G_{f}=\mathbb{Z}_{2}^{f}\times G_{b}$, where $G_{b}$ is the group extension of $\mathbb{Z}_{2}^{\mathcal{T}}$ by $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ with $\rho_{\mathcal{T}}$ and $\nu_{2}^{b}$. In total, there are twelve groups with $\nu_{2}^{b}=1$ (Table 1) and five groups with $\nu_{2}^{b}=\mathbf{g}^{2m}$ (Table 2).
@@
Conclusion and discussion — In summary, we have demonstrated that $\mathcal{T}^{2}=1$ TRS is enforced to break by a kind of exotic 3D fTO, featuring Ising non-Abelian three-loop braiding statistics. However, the usual TRS with $\mathcal{T}^{2}=\mathcal{P}_{f}$ is still compatible with such a 3D fTO. We have also discussed a framework for exploring ESB for generic 3D fTOs. We further conjecture that parity violation in the Standard Model might arise from the vacuum being certain Ising 3D fTO, and it is a kind of ESB phenomenon.

### v2

@@ Abstract
In the previous work Ref.[13], we systematically studied the ESB for 2D invertible fTO. Herein, we venture into the ESB phenomenon of 3D fTO. As there is no invertible fTO in 3D, we have to consider intrinsic fTO. In this work, we will focus on 3D ESB of TRS. We show that TRS is intrinsically broken by the presence of certain 3D fermionic TO (fTO) characterized by exotic non-Abelian three-loop braiding statistics. Such an fTO is closely related to an intrinsic interacting fermionic SPT (fSPT) order, which is specifically protected by symmetry $K_{f}=\mathbb{Z}_{2}^{f}\times\mathbb{Z}_{2}\times\mathbb{Z}_{8}$, characterized by non-Abelian three-loop braiding statistics [14]. This fTO is obtained by gauging $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ in the fSPT state, and will be referred to as Ising fTO. As we shall show, this $K_{f}$ fSPT is incompatible with $\mathcal{T}^{2}=1$ while compatible with $\mathcal{T}^{2}=\mathcal{P}_{f}$. Its incompatibility with the $\mathcal{T}^{2}=1$ TRS becomes exactly the ESB of TRS by this fTO. To our knowledge, this is the first and simplest example of TRS incompatibility of 3D fTO. This example readily extends to the realm of parity symmetry through the crystalline equivalence principle, while the concept of ESB unveils a novel pathway for comprehending parity violation within the Standard Model (SM), under the assumption that its vacuum acquires emergent generalized symmetry[15, 16, 17] characterized as certain 3D Ising fTO. We will also discuss the general framework for understanding 3D ESB phenomena.
@@
Upon gauging $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ in this fSPT (promoting the global $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ symmetry to be a local gauge symmetry), we obtain a 3D fTO, dubbed Ising fTO, with two elementary particle-like topological excitations: gauge charges $\mathsf{e}_{1,2}$, together with two elementary loop excitations $\Omega_{1,2}$, containing unit flux of $\mathbb{Z}_{2}$ and $\mathbb{Z}_{8}$ gauge group, respectively. Besides, there is also fermionic excitation that can be created by local operators and is usually called the transparent fermion $f$. General loop excitations contain the gauge flux and gauge charge. The flux of loop excitations forms the fusion group $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$. This fTO has many exotic properties. First, when $\Omega_{1}$ is linked to the base loop $\Omega_{2}$, its topological spin $\theta_{\Omega_{1};\Omega_{2}}$ satisfies $\Theta_{\Omega_{1};\leavevmode\nobreak\ \Omega_{2}}=2\theta_{\Omega_{1};\leavevmode\nobreak\ \Omega_{2}}=\frac{\pi}{4}$ mod $\frac{\pi}{2}$. Secondly, for two linked $\Omega_{1}$ and $\Omega_{2}$, each carries a Majorana zero mode (MZM), indicating the non-Abelian nature (see Fig.1c) [19, 14]. Thirdly, when further gauging the fermion parity, this resulting TO has fascinating statistics among loops, one of which is called the non-Abelian Ising three-loop braiding statistics as seen in Fig.1b . We note that in general the topological invariant $\Theta_{\Omega_{1};\leavevmode\nobreak\ \Omega_{2}}=\frac{x\pi}{4}$ mod $2\pi$, and it is non-Abelian when $x$ takes odd integer. In addition, there is another topological invariant $\Theta_{\Omega_{2};\leavevmode\nobreak\ \Omega_{1}}$ which can take 00 or $\pi$ characterizing an Abelian bosonic TO of $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ Dijkgraaf-Witten models. (For completeness, a concise overview of fSPT and associated three-loop braiding is provided in the supplementary materials). We also note that the ESB by TRS or parity symmetry we discuss below applies to these fTOs for all odd $x$.
@@
For three-loop statistics, this requirement leads to the conditions $\Theta_{\Omega_{1};\Omega_{2}}=-\Theta_{\rho(\Omega_{1});\leavevmode\nobreak\ \rho(\Omega_{2})}$ and $\Theta_{\Omega_{2};\Omega_{1}}=-\Theta_{\rho(\Omega_{2});\leavevmode\nobreak\ \rho(\Omega_{1})}$, where
@@
(see End Matter for derivations). Solving the equations, we find (i) when $t=0$, $p=3$; (ii) when $t=1$, $p=3$ if $\Theta_{\Omega_{1},\Omega_{2};\leavevmode\nobreak\ \Omega_{2}}=0$ and $p=1$ if $\Theta_{\Omega_{1},\Omega_{2};\leavevmode\nobreak\ \Omega_{2}}=\pi$; (iii) $q=0$. Therefore, we are left with only three possibly consistent $\rho$, labeled by $\mathbf{r}=(3,0,0)$, $(3,0,1)$, and $(1,0,1)$, respectively.
@@
The structure of $G_{f}$ is described by: (1) how TRS acts on the subgroup $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$: $\rho_{\mathcal{T}}(\mathbf{g})=\mathbf{g}^{1+2p}\mathbf{h}^{q}$ and $\rho_{\mathcal{T}}(\mathbf{h})=\mathbf{g}^{4t}\mathbf{h}$, where $\mathbf{r}=(p,q,t)$ is the same as above [31]; and (2) what $\mathcal{T}^{2}$ is: $\mathcal{T}^{2}=\nu_{2}^{a}\nu_{2}^{b}\in K_{f}$, where $\nu_{2}^{a}=1$ or $\mathcal{P}_{f}$, and $\nu_{2}^{b}=1,\mathbf{g}^{2},\mathbf{g}^{4}$ for $\mathbf{r}=(3,0,0)$ or $(p,1,0)$ and $\nu_{2}^{b}=1$ for other choices of $\mathbf{r}$ (after considering some physical constraints for simplification  [^2] There are in total 28 different $G_{b}$, which will simplify to 17 cases after these physical constraints. However, we calculate all the 28 cases and found no case can reduce to the $K_{f}$ fSPT. ; see End Matter). All 12 choices of $\mathbf{r}$ are considered, not limited to the three choices compatible with three-loop statistics. The Kramer-singlet and -doublet fermion systems correspond to $\nu_{2}^{a}=1$ and $P_{f}$, respectively. Below we demonstrate that ESB occurs for Kramer-singlet fermions. In this case, $G_{f}=\mathbb{Z}_{2}^{f}\times G_{b}$, where $G_{b}$ is the group extension of $\mathbb{Z}_{2}^{\mathcal{T}}$ by $\mathbb{Z}_{2}\times\mathbb{Z}_{8}$ with $\rho_{\mathcal{T}}$ and $\nu_{2}^{b}$. In total, there are twelve groups with $\nu_{2}^{b}=1$ (Table 1) and five groups with $\nu_{2}^{b}=\mathbf{g}^{2m}$ (Table 2).
@@
Conclusion and discussion — In summary, we demonstrate that $\mathcal{T}^{2}=1$ TRS is enforced to break by a kind of exotic 3D fTO, featuring Ising non-Abelian three-loop braiding statistics. However, the usual TRS with $\mathcal{T}^{2}=\mathcal{P}_{f}$ is still compatible with such a 3D fTO. We also discussed a framework for exploring ESB for generic 3D fTOs. We further conjecture that parity violation in the Standard Model might arise from the vacuum being certain Ising 3D fTO, and it is a kind of ESB phenomenon.